OpenStudy (jiteshmeghwal9):

Find general solution of \(x\) in the equation :- \[sin6x=sin4x-sin2x\]

OpenStudy (michele_laino):

hint: your equation, can be rewritten as follows: \[\Large \sin \left( {4x} \right) = \sin \left( {6x} \right) + \sin \left( {2x} \right)\]

OpenStudy (jiteshmeghwal9):

\[\sin4x=2\sin4xcos2x\]\[\sin4x(1-2\cos2x)=0\]\[x=\frac{n \pi}{4}\]&\[2x=2n \pi \pm \frac{\pi}{3}\]\[x=n \pi \pm \frac{\pi}{6}\]

OpenStudy (michele_laino):

yes! exact that was my idea

OpenStudy (jiteshmeghwal9):

thanx

OpenStudy (michele_laino):

:)